# Symmetries of Love: Ladder Structure of Static and Rotating Black Holes

### 02 Jan 2024 v2

@@ I Introduction
where $\Delta_{b}(r)=r^{2}-c_{2}\,r+c_{3}$ with $(c_{2},c_{3})$ being some constants and $h(r)$ is an arbitrary radial function. We also extend our analysis to find the most general metric within the so-called Konoplya-Rezzolla-Zhidenko class that has a similar ladder structure.
It is intriguing that the imposition of the ladder symmetry leads to such severe constraints on the form of the BH metric. Moreover, because of ladder symmetry, these BHs will have vanishing TLN for scalar perturbations. Considering that astrophysical BHs are rarely isolated and are under constant external influence, our explorations may enhance the understanding of how BHs behave in the presence of perturbations—an aspect of central observational importance Fang:2005qq ; Creci:2021rkz ; Cai:2019npx ; Tan:2020hog ; Charalambous:2023jgq ; Katagiri:2023umb ; DeLuca:2022xlz ; Katagiri:2023yzm ; Zi:2023pvl ; Nair:2022xfm ; Chakraborty:2023zed .
@@ IV Conclusion and Discussions
Motivated by this important result, we have presented the most general static and stationary (in KRZ class) BH metrics having ladder symmetry for static (with frequency $\omega=0$) scalar perturbations. This in turn implies that the corresponding BHs have zero TLN for $\omega=0$, i.e., $\Lambda=\mathcal{O}(M\,\omega)$. Actually, for the case of static BHs, our result is even stronger. In particular, since $H_{\ell}\,\phi_{\ell}\propto\omega^{2}$ for non-static perturbations, the corresponding TLN must be $\mathcal{O}(M^{2}\omega^{2})$. However, a similar assertion does not hold for the rotating case, because $H_{\ell}\,\phi_{\ell}\propto\omega$ for non-static perturbations. These conclusions match with the recent claim reported in Ref. Perry:2023wmm .
Given the immense theoretical and observational significance, it will be interesting to extend our work for gravitational perturbations. Another important prospect would be studying various properties like geodesic structure, shadow, and stability of the general class of BH spacetimes given by Eq. (19) and Eq. (20). We leave these analyses for a future attempt.
@@ V Acknowledgement
CS acknowledges the support from the Sabarmati Bridge Fellowship (Project ID: MIS/IITGN-SBF/PHY/SS/2023-24/025) from IIT Gandhinagar. The research of R.G. is supported by the Prime Minister Research Fellowship (PMRF ID: 1700531), Government of India. S. S. acknowledges support from the Department of Science and Technology, Government of India under the SERB CRG Grant (CRG/2020/004562).

### 23 Feb 2024 v3

# Exploring Ladder Symmetry and Love Numbers for Static and Rotating Black Holes

@@ I Introduction
where $\Delta_{b}(r)=r^{2}-c_{2}\,r+c_{3}$ with $(c_{2},c_{3})$ being some constants and $h(r)$ is an arbitrary radial function. The zeros of the function $\Delta_{b}(r)$ determine the location of the horizon. We also extend our analysis to find the most general metric within the so-called Konoplya-Rezzolla-Zhidenko class that has a similar ladder structure.
It is intriguing that the imposition of the ladder symmetry leads to such severe constraints on the form of the BH metric. The same symmetry can then be used to conclude the vanishing of the tidal Love number (TLN) associated with these black hole solutions. Though the existence of the ladder symmetry is not limited to black holes alone, it is important to note that the mere presence of ladder symmetry does not automatically ensure the vanishing of the TLN. In fact, to establish that TLN vanishes from ladder symmetry requires additionally the presence of a regular horizon. Specifically, this condition is crucial for demonstrating that the zeroth mode ($\ell=0$) of the perturbation have a zero TLN. Only then can one show that all higher modes also have zero TLN by a repeated application of the raising operator. Considering that astrophysical BHs are rarely isolated and are under constant external influence, our explorations may enhance the understanding of how BHs behave in the presence of perturbations—an aspect of central observational importance Fang:2005qq ; DeLuca:2021ite ; Creci:2021rkz ; Cai:2019npx ; Tan:2020hog ; Charalambous:2023jgq ; Katagiri:2023umb ; DeLuca:2022xlz ; Katagiri:2023yzm ; Zi:2023pvl ; Nair:2022xfm ; Chakraborty:2023zed .
@@ IV Conclusion and Discussions
Motivated by this important result, we have presented the most general static and stationary (in KRZ class) BH metrics having ladder symmetry for static (with frequency $\omega=0$) scalar perturbations. This in turn implies that the corresponding BHs have zero TLN for $\omega=0$, i.e., $\Lambda=\mathcal{O}(M\,\omega)$. Actually, for the case of static BHs, our result is even stronger. In particular, since $H_{\ell}\,\phi_{\ell}\propto\omega^{2}$ for non-static perturbations, the corresponding TLN must be $\mathcal{O}(M^{2}\omega^{2})$. However, a similar assertion does not hold for the rotating case, because $H_{\ell}\,\phi_{\ell}\propto\omega$ for non-static perturbations. These conclusions match with the recent claim reported in Ref. Perry:2023wmm . We have also found that for these background spacetimes, the ladder symmetry is also a symmetry at the level of the action of these perturbing fields.
Given the immense theoretical and observational significance, it will be interesting to extend our work for gravitational perturbations as well as for non-liner perturbations DeLuca:2023mio . Another important prospect would be studying various properties like geodesic structure, shadow, and stability of the general class of BH spacetimes given by Eq. (19) and Eq. (20). We leave these analyses for a future attempt.
@@ V Acknowledgement
We thank the anonymous referee for useful comments on our work. CS acknowledges the support from the Sabarmati Bridge Fellowship (Project ID: MIS/IITGN-SBF/PHY/SS/2023-24/025) from IIT Gandhinagar. The research of R.G. is supported by the Prime Minister Research Fellowship (PMRF ID: 1700531), Government of India. S. S. acknowledges support from the Department of Science and Technology, Government of India under the SERB CRG Grant (CRG/2020/004562).